b)  Answer the following stoichiometry questions by referring to the equation below:

2        KClO3         ---->      2 KCl + 3 O2

i.        If 1.50 mol of KClO3 decomposes, what is the mass of O2 that will be produced?

ii.       If 80.0 grams of O2 was produced, how many moles of KClO3  are decomposed?

iii.    Find the mass of KClO3 needed if we need to produce 2.75 mol of KCl.

1
Expert's answer
2021-03-08T07:09:17-0500

i)moles of O2 produced are 4.5 moles

Rmm = 32

Mass=32 × 4.5= 144g

ii) "Moles= mass\/rmm"


=122.6 /80

=1.53moles

iii) moles of 2.75/2= 1.375moles

Mass= 1.375× 122.6=168.6g